| Operation Summary for December 9, 2011 - December 12, 2011 |
|
NuMI reported receiving 9.58E18 protons on target for the period from 12/5/11 to 12/12/11. Friday December 9 The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, and with the MiniBooNE and NuMI experiments taking beam. During the day shift, an Instrumentation person reported that he had repaired the MI-40 BPM. Booster personnel began a study at 9:25 AM. Muon Rings personnel began taking 30 minutes of beam at 10:17 AM to help an Instrumentation person diagnose the P1 transfer line BPM problem. Booster studies ended at 12:37 PM. At 1 PM, operations halted beam for about an hour for work on the MI RF system, the Booster Chopper, on a Booster Correction Element power supply (B:CEPS400), and to add water to the MI LCW system. At 2:50 PM, B:CEPS400 tripped off due to a seized fan; Booster personnel investigated and reset it. At 6:38 PM, several vacuum pumps in the P1 transfer line tripped off; operators reset them. During the evening shift, Operations established beam to the Meson FTBF experiments T-1015 and T-1017 at 5:38 PM. B:CEPS400 tripped off at 9:11 PM; it reset. Operations disabled T-1015 and T-1017 beam at 10:01 PM. Saturday December 10 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 4:17 AM, operators reported a power outage for MTA. Operators contacted the Cryo Coordinator; a yard breaker had tripped. The Duty Electrician restored the power and Cryo personnel brought back the MTA cryo system. During the day shift, Operations established beam to T-1015 and T-1017 at 10:17 PM. At 11:39 AM, operators reported that the MI-10 Power Supply Interlock Loop (PSLOOP) tripped off and they had lost communication with MI-10; operators investigated and found the PLC not working. Operators contacted Electrical Support personnel and put in an energy conservation time line into the Time Line Generator (TLG). MI beam returned at 2:18 PM. EE Support personnel reported that they had replaced a CPU. During the evening shift, at 5:41 PM, a Booster vacuum pump (B:IPRF22) tripped off and caused two Booster RF stations (BRF11 and 12) to also trip off; everything reset. At 7:50 PM, BRF11 tripped off; it reset. At 9:36 PM, the NuMI beam permit tripped off when a kicker (I:KPS6N) didn’t fire; everything reset. Operations disabled T-1015 and T-1017 beam at 10 PM. Sunday December 11 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 12:46 AM, Linac RF station (LRF4) tripped off for the first of two times in two hours; it reset. At 1:23 AM, LRF2 tripped off for the first of two times in two minutes; it reset. At 2:54 AM, an I- Source vacuum pump tripped off; it reset. At 6:09 AM, BRF11 tripped off; it reset. At 6:48 AM, a Linac quadrupole magnet power supply (L:QPS107) alarmed on a low output; operators tried adjusting it, but ended up swapping out its control cards. During the day shift, at 8:52 AM, MIRF12 tripped off for the first of three times in about 30 minutes; it reset. Operations established beam to T-1015 and T-1017 at 10:49 AM. During the evening shift, at 6:05 PM, operators reported a vacuum burst in the P1 transfer line; operators reset vacuum pumps. Operators completed a Booster injection tune at 7:29 PM. The NuMI beam permit tripped off at 8:28 PM on a MI RF gap indication; it reset. Operations disabled T-1015 and T-1017 beam at 10:06 PM; experiment T-1017 competed its run. Operators completed a Booster tune injection at 11:53 PM. Monday December 12 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, problems with a Linac quadrupole power supply (L:QPS107) came up again; operators turned off the Linac beam for about 15 minutes while they swapped out control cards. Operations Coordinator (OpsCo) The OpsCo approved of the plans by Booster, Muon, and MI personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 12/12/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |